OpenStudy (anonymous):

Line m has slope -4 and contains the point (-6, -3). Which of the following is a point-slope equation of line m?

OpenStudy (anonymous):

y+3 = -4(x+6)

Directrix (directrix):

We'll use the point-slope form for a line. y - y1 = m(x - x1) where (x1,y1) is (-6, -3) and m = -4 y - (-3) = (-4) ( x - (-6)) y + 3 = -4(x +6) y + 3 = -4x -24 y = -4x -27 ----check my work --->